# New simple and accurate approximations for quintessence

### v2

@@ Abstract
We derive new simple approximations for quintessence solutions to Einsteins’s field equations in spatially homogeneous, isotropic and flat spacetimes. The approximations are typically an order of magnitude more accurate than anything available in the literature, which from an observational perspective makes numerical calculations superfluous. For example, our tracking quintessence approximation yields a $\sim 0.1\%$ maximum relative error of $H(z)/H_{0}$ for the observationally viable inverse power law scalar field potentials, and similarly for viable thawing quintessence models using two slow-roll parameters. We situate the new results in a broader dark energy context, where the present parametrized deviations from $\Lambda$CDM offer simple alternatives to the CPL parametrization, useful for observations at increasingly large redshifts.
@@ 1 Introduction
A growing number of increasingly precise cosmological observations of, e.g., type Ia supernovae [1, 2, 3], the cosmic microwave background (CMB) [4], various measurements of the Hubble parameter [5, 6, 7], the universe large-scale structure (LSS) [8, 9], and baryon acoustic oscillations (BAO) [10, 11, 12, 13], suggests that the universe is presently accelerating. In addition cosmological observations indicate that most of the matter in the universe is invisible ‘dark matter’ [14, 15, 16, 17]. The standard interpretation of the observations rely on that General Relativity is an accurate description of gravity on cosmological scales and that the observable Universe is almost spatially homogeneous, isotropic and flat on sufficiently large spatial scales. This leads to that the acceleration of the universe requires a ‘dark energy’ (DE) density component, $\rho_{\mathrm{DE}}$, with sufficiently large negative pressure, while the dark matter seems to be best described by a dominant cold dark matter (CDM) energy density component, $\rho_{\mathrm{m}}$, without pressure. The most simple and remarkably successful model that is compatible with these assumptions is the $\Lambda$CDM model, for which the energy density $\rho_{\mathrm{DE}}$ and pressure $p_{\mathrm{DE}}$ are given by $\rho_{\mathrm{DE}}=-p_{\mathrm{DE}}=\Lambda>0$, where $\Lambda$ is the cosmological constant.
@@
There are, however, theoretical problems, e.g., the nature of dark matter [18, 19], the observed small value of $\Lambda$, the coincidence problem, see e.g. [20] and references therein. Recent tensions between some data sets [21, 22, 23] and that large scale structures seem to form surprisingly early in the universe [24] have further aggravated the situation, which has contributed as motivation for numerous dynamical DE models for which $\rho_{\mathrm{DE}}$ is evolving. Many of these models are purely cosmographically motivated, without any tie to any theory, notably the Chevallier-Polarski-Linder (CPL) parametrization [25, 26] of the DE equation of state parameter $w_{\mathrm{DE}}=p_{\mathrm{DE}}/\rho_{\mathrm{DE}}$ and a plethora of variations of it, see e.g. [27] and references therein.
@@
Some DE models, however, endeavour to have some theoretical motivation, where one of the simplest attempts is quintessence, i.e. a canonical scalar field, $\phi$, that is minimally coupled to gravity, see e.g. [28] and [29] for references. This is the main topic of the present paper, which deals with a non-interacting scalar field and matter in a spatially homogeneous, isotropic and flat spacetime, although note that the ideas in this paper are applicable to many other models and alternative theories. More specifically, we will derive simple and accurate approximations for thawing and tracking quintessence solutions, [^1] According to Tsujikawa (2013) [28] there are three types of quintessence in the literature: thawing, tracking and scaling freezing quintessence, see [30] for subtleties. In this paper we consider tracking quintessence for asymptotically inverse power law scalar field potentials $V\propto\phi^{-p}$, $p>0$, when $\phi\rightarrow 0$. Since scaling freezing quintessence does not yield past continuous deformations of $\Lambda$CDM we will only briefly touch on this topic in appendix B. which yield continuous deformations of $\Lambda$CDM cosmology. Notably, the approximations are typically simpler and an order of magnitude more accurate than anything available in the literature. For example, our tracking quintessence approximation yields a $\sim 0.1\%$ maximum relative error of $H(z)/H_{0}$ for the observationally viable inverse power law scalar field potentials, and similarly for viable thawing quintessence models using two slow-roll parameters. We note that from an observational perspective the remarkable accuracy of our approximations makes numerical calculations superfluous when assessing observational constraints; moreover the approximations yield an analytic overview of the observational viability of broader ranges of models than those that are typically numerically investigated. For example, it would be interesting to replace the recent numerical observational constraint analysis by Shlivko and Steinhardt [31] with our thawing quintessence approximations for a broader set of models, although we note that our new parametrized DE deviations from $\Lambda$CDM offer simple alternatives to the CPL parametrization, useful for observations at increasingly large redshifts.
@@
The outline of the paper is as follows. Motivated by our new quintessence approximations, the next section provides an associated general DE context. In Section 3 we derive the expressions for the parameters appearing in the previous DE section for thawing and tracking quintessence. The new thawing and tracking quintessence approximations are compared with numerical calculations and previous results in the literature in section 4. The main part of the paper ends with section 5, which includes suggestions for future applications of the present approach and results. Appendix A reviews quintessence approximations found in the literature, while appendix B deals with scaling freezing quintessence.
2 Dark energy considerations
@@ 2.1 Preliminaries
Since the spacetime is described by the Hubble parameter $H(z)$ it follows that cosmological observables and their interpretation depend directly on $H(z)$, its derivatives and integrals. This is illustrated by, e.g., the deceleration parameter $q=-1-\dot{H}/H^{2}=-1+(1+z)\frac{dH}{dz}/H$, and many distance measures, e.g., the Hubble distance $D_{H}(z)$, the proper motion or coordinate distance $D_{M}(z)$, the luminosity distance $D_{L}(z)$, the angular diameter distance $D_{A}(z)$, which are given by (see, e.g. [27])
@@ 2.2 Relationships between $\Omega_{\mathrm{DE}}$, $\Omega_{\mathrm{m}}$, $w_{\mathrm{DE}}$ and $H$
Next we introduce the dimensionless and bounded Hubble normalized variables $\Omega_{\mathrm{DE}}$ and $\Omega_{\mathrm{m}}$:
@@
which yields [^4] To improve our approximations we will use $\Omega_{\mathrm{DE}}(N)$, but subsequently the relation $\Omega_{\mathrm{DE}0}=1-\Omega_{\mathrm{m}0}=1-\rho_{\mathrm{m}0}/3H_{0}^{2}$ can be used to bring the approximations closer to present day data.
@@
If one instead of $w_{\mathrm{DE}}(N)$ has $\Omega_{\mathrm{DE}}(N)$, then one can obtain $w_{\mathrm{DE}}(N)$ and $E_{\mathrm{DE}}^{2}(N)$ since
@@
$E_{\mathrm{DE}}^{2}$ $=\exp(-3N)\left(\frac{\Omega_{\mathrm{m}0}}{\Omega_{\mathrm{m}}}\right)=\exp(-3N)\left(\frac{1-\Omega_{\mathrm{DE}0}}{1-\Omega_{\mathrm{DE}}}\right),$ (14b)
@@
$q=-\frac{\ddot{a}}{aH^{2}}=-1-\frac{H^{\prime}}{H},$ (15)
@@
In the following sections we will discuss quintessence for which $w_{\mathrm{DE}}=w_{\phi}$. As shown in section 3, see also e.g. [30] and [32], thawing quintessence yields $\lim_{N\rightarrow-\infty}w_{\phi}=-1$ while tracking quintessence leads to $\lim_{N\rightarrow-\infty}w_{\phi}=w_{\infty}=\mathrm{constant}$, $-1<w_{\infty}<0$. As a consequence thawing and tracking quintessence are asymptotically toward the past described by the $\Lambda$CDM and $w$CDM models, respectively. Let us therefore recall the properties of these models.
@@ 2.4 Past DE series expansions
Thawing and tracking quintessence, as illustrated in a dynamical systems setting in [30] and [32], are associated with attractor solutions that originate from fixed points on the matter dominant boundary where $\Omega_{\phi}=0$ and $\Omega_{\mathrm{m}}=1$. Treating both cases simultaneously as dark energy models, i.e., by letting $w_{\mathrm{DE}}\rightarrow w_{\infty}\in[-1,0)$ and $\Omega_{\mathrm{DE}}\rightarrow 0$ when $N\rightarrow-\infty$, leads to that equation (14a), which can be written as $\Omega_{\mathrm{DE}}^{\prime}=-3w_{\mathrm{DE}}\Omega_{\mathrm{DE}}(1-\Omega_{\mathrm{DE}})$, results in the approximation $\Omega_{\mathrm{DE}}^{\prime}=-3w_{\infty}\Omega_{\mathrm{DE}}$ in this limit and hence
$\Omega_{\mathrm{DE}}\approx Ce^{-3w_{\infty}N},$ (20)
which is indeed the case for the thawing ($w_{\infty}=-1$) and tracking ($w_{\infty}\in(-1,0)$) quintessence models. To improve the approximation we now perform series expansions in
@@
where we for brevity only consider $w_{\mathrm{DE}}$ and $\Omega_{\mathrm{DE}}$ (note that $T\rightarrow 0$ when $N\rightarrow-\infty$). If one starts with a series expansion in $w_{\mathrm{DE}}$ and uses (7) and (13) one obtains an expression for $\Omega_{\mathrm{DE}}$ that subsequently can be series expanded in $T$; on the other hand, if one starts with a series expansion in $\Omega_{\mathrm{DE}}$ and inserts this into (14a) one obtains an expression for $w_{\mathrm{DE}}$ that one then can series expand in $T$. These two procedures result in the same series expansions, which can be written as [^5] The given expressions can be generalized by writing $w_{\mathrm{DE}}=w_{\infty}\left[1+\alpha T+\sigma T^{2}+\dots\right]$, which subsequently can be used to compute the series expansions for $\Omega_{\mathrm{DE}}$ and, e.g., $E_{\mathrm{DE}}^{2}$. The reason we have written $\alpha$ as $1-\gamma$ and $\sigma$ as $(\gamma-1)\beta$ is that this simplifies and is compatible with our thawing and tracking quintessence approximations, where these use the $[1/1]$ Padé expression for $\Omega_{\mathrm{DE}}$ in order to connect our approximations to the present time.
@@ 2.5 DE Padé approximants
For brevity we will restrict our considerations by following the most common approach in the literature, which is to use $w_{\mathrm{DE}}$ as the starting point. The series expansions provide approximations when $T$ is small, i.e. close to the matter dominated regime. To obtain approximations that are accurate even up to the present time we will use Padé approximants for the series expansion of $w_{\mathrm{DE}}$ in $T$. Our first and second order expansions in $T$ for $w_{\mathrm{DE}}$ yield the following Padé approximants for $w_{\mathrm{DE}}$:
$w_{\mathrm{DE}}$ $\approx[0/1]_{w_{\mathrm{DE}}}(T)=\frac{w_{\infty}}{1+(\gamma-1)T},$ (23a)
$w_{\mathrm{DE}}$ $\approx[1/1]_{w_{\mathrm{DE}}}(T)=w_{\infty}\left(1-\frac{(\gamma-1)T}{1+\beta T}\right).$ (23b)
@@
To connect with the present time we can set $N=0$ and solve for $T=T_{0}$ in (23) in terms of the left hand side $w_{\mathrm{DE}0}$. However, since $w_{\mathrm{DE}0}$ is not an observable we will instead relate $T_{0}$ to $\Omega_{\mathrm{DE}0}$ and thereby obtain a direct connection to the observables $H_{0}$ and $\rho_{\mathrm{m}0}$ since $1-\Omega_{\mathrm{DE}0}=\Omega_{\mathrm{m}0}=\rho_{\mathrm{m}0}/3H_{0}^{2}$. To obtain accurate approximations at the present time for $\Omega_{\mathrm{DE}}$ we consider the Padé approximant (use the expression (22b) up to second order in $T$)
$\Omega_{\mathrm{DE}}\approx[1/1]_{\Omega_{\mathrm{DE}}}(T)=\frac{T}{1+\gamma T},$ (24)
@@
from which it follows that $\Omega_{\mathrm{DE}0}=T_{0}/[1+\gamma T_{0}]$, which yields

### v3

# New simple and accurate quintessence approximations

@@ Abstract
We derive new approximations for quintessence solutions that are simpler and an order of magnitude more accurate than anything available in the literature, which from an observational perspective makes numerical calculations superfluous. For example, our tracking quintessence approximation yields $\sim 0.1\%$ maximum relative errors of $H(z)/H_{0}$ and $\Omega_{\mathrm{m}}(z)$ for the observationally viable inverse power law scalar field potentials, and similarly for viable thawing quintessence models using two slow-roll parameters. The approximations are trivially computed from the scalar field potential and as an application we give analytic expressions for the CPL parameters calculated from an arbitrary scalar field potential for thawing and tracking quintessence models.
@@ 1 Introduction
A growing number of increasingly precise cosmological observations of, e.g., type Ia supernovae [1, 2, 3], the cosmic microwave background (CMB) [4], various measurements of the Hubble parameter [5, 6, 7], the universe large-scale structure (LSS) and baryon acoustic oscillations (BAO) [8, 9, 10, 11, 12, 13, 14, 15], suggest that the universe is presently accelerating. In addition cosmological observations indicate that most of the matter in the universe is invisible ‘dark matter’ [16, 17, 18, 19]. The standard interpretation of the observations rely on that General Relativity is an accurate description of gravity on cosmological scales and that the observable Universe is almost spatially homogeneous, isotropic and flat on sufficiently large spatial scales. This leads to that the acceleration of the universe requires a ‘dark energy’ (DE) density component, $\rho_{\mathrm{DE}}$, with sufficiently large negative pressure, while the dark matter seems to be best described by a dominant cold dark matter (CDM) energy density component, $\rho_{\mathrm{m}}$, without pressure. The simplest model that is compatible with these assumptions is the $\Lambda$CDM model, for which the energy density $\rho_{\mathrm{DE}}$ and pressure $p_{\mathrm{DE}}$ are given by $\rho_{\mathrm{DE}}=-p_{\mathrm{DE}}=\Lambda>0$, where $\Lambda$ is the cosmological constant.
@@
There are, however, theoretical problems, e.g., the nature of dark matter [20, 21], the observed small value of $\Lambda$ and the coincidence problem, see e.g. [22] and references therein. Recent tensions between some data sets [23, 24, 25] and that large scale structures seem to form surprisingly early in the universe [26] have further aggravated the situation, which has contributed as motivation for numerous dynamical DE models for which $\rho_{\mathrm{DE}}$ is evolving. Many of these models are purely cosmographically motivated, without any tie to any theory, notably the Chevallier-Polarski-Linder (CPL) parametrization [27, 28] of the DE equation of state parameter $w_{\mathrm{DE}}=p_{\mathrm{DE}}/\rho_{\mathrm{DE}}$ and a plethora of variations thereof, see e.g. [29] and references therein.
@@
Some DE models, however, endeavour to have some theoretical motivation, where one of the simplest attempts is quintessence, i.e. a canonical scalar field, $\phi$, minimally coupled to gravity, see e.g. [30] and [31] for references. This is the topic of the present paper, which deals with a non-interacting scalar field and matter in a spatially homogeneous, isotropic and flat spacetime, although note that the ideas in this paper are applicable to many other models and alternative theories. More specifically, we will derive simple, analytic, and accurate approximations for thawing quintessence, which exists for all scalar field potentials, and tracking quintessence for asymptotically inverse power-law potentials, i.e. $V\propto\phi^{-p}$, $p>0$ when $\phi\rightarrow 0$ (i.e. the types of quintessence that dominate the observational literature). [^1] Different types and nomenclatures for quintessence are discussed in section 3. Notably, our best approximations are typically simpler and an order of magnitude more accurate than anything available in the literature. For example, our tracking quintessence approximation yields $\sim 0.1\%$ maximum relative errors for $H(z)/H_{0}$ and $\Omega_{\mathrm{m}}(z)$ for observationally viable inverse power law scalar field potentials, and similarly for viable thawing quintessence models using two slow-roll parameters. We note that from an observational perspective the remarkable accuracy of our approximations makes numerical calculations superfluous when assessing observational constraints; moreover the approximations yield an analytic overview of the observational viability of broader ranges of models than those that are typically numerically investigated.
The accuracy and scope of our approximations have far ranging consequences. There are many papers that start from specific scalar field potentials and numerically compute observational consequences for a variety of initial data, often not particularly numerically systematically explored. For a wide range of scalar field potentials the present paper makes numerical explorations mote. Although there are mathematical considerations in this paper that might be unfamiliar to people in this field, the results are very easy to understand. For the reader not interested in the derivation of the approximations and only wants to use the results, just plug in your favourite potential into our master equations: for thawing quintessence use eq. (47) to obtain approximations for $w_{\phi}(z)$, $\Omega_{\phi}(z)$ (or, equivalently, $\Omega_{\mathrm{m}}(z)=1-\Omega_{\phi}(z)$) and $H(z)$, where $z$ is the redshift, while eq. (69) yields analytic expressions for the CPL parameters, thereby obtaining simple, accurate, analytical descriptions of the recent analysis of Wolf and Ferreira [32] and by Shlivko and Steinhardt [33], where one can also broaden the range of models; for tracking quintessence use eqs. (57) and (58) to obtain approximations for $w_{\phi}(z)$, $\Omega_{\phi}(z)$ (or, equivalently, $\Omega_{\mathrm{m}}(z)=1-\Omega_{\phi}(z)$) and $H(z)$, while eq. (69) gives analytic expressions for the CPL parameters.
@@
The outline of the paper is as follows. The next section provides a unifying general DE context for some of the structures that thawing and tracking quintessence share. In Section 3 we derive the expressions for the parameters appearing in the previous DE section that are particular for thawing and tracking quintessence, respectively. In addition, we explicitly give the simple analytical expressions for $w_{\phi}(z)$, $\Omega_{\phi}(z)$ (or, equivalently $\Omega_{\mathrm{m}}(z)=1-\Omega_{\phi}(z)$) and $H(z)$. Section 4 gives explicit analytical expressions for the CPL parameters based on our new extremely accurate approximations. From an observational standpoint, this eliminates the need for numerical calculations of the CPL parameters for whatever potential one might be interested in, which is how one traditionally computes them for quintessence, see e.g. [32, 33]. Section 5 compares the new thawing and tracking quintessence approximations with numerical calculations and previous approximations in the literature, reviewed in Appendix A. The main part of the paper ends with section 6, which includes suggestions for future applications of the present approach and results. Appendix B deals with scaling freezing quintessence (the remaining type of quintessence not covered in the main part of the paper since $\lim_{z\rightarrow\infty}\Omega_{\phi}>0$ in this case, which is in contrast to thawing and tracking quintessence for which $\lim_{z\rightarrow\infty}\Omega_{\phi}=0$).
2 Common thawing and tracking quintessence considerations
In this section we will consider rather general aspects of dark energy that cover some features that are common for both thawing and tracking quintessence, which we treat individually in section 3.
@@ 2.1 Preliminaries
Since the spacetime is described by the Hubble parameter $H(z)$ it follows that cosmological observables and their interpretation depend directly on $H(z)$, its derivatives and integrals. This is illustrated by, e.g., the deceleration parameter $q=-a\ddot{a}/\dot{a}^{2}=-1+(1+z)\frac{dH}{dz}/H$, and many distance measures, e.g., the Hubble distance $D_{H}(z)$, the proper motion or coordinate distance $D_{M}(z)$, the luminosity distance $D_{L}(z)$, the angular diameter distance $D_{A}(z)$, which are given by (see, e.g. [29])
@@ 2.2 Relationships between $\Omega_{\mathrm{DE}}$, $\Omega_{\mathrm{m}}$, $w_{\mathrm{DE}}$ and $H$
Next we introduce the dimensionless and bounded Hubble-normalized variables $\Omega_{\mathrm{DE}}$ and $\Omega_{\mathrm{m}}$:
@@
which yields
@@
If one instead of $w_{\mathrm{DE}}(N)$ has $\Omega_{\mathrm{DE}}(N)$ (or $\Omega_{\mathrm{m}}(N)=1-\Omega_{\mathrm{DE}}(N)$), then one can obtain $w_{\mathrm{DE}}(N)$ and $E_{\mathrm{DE}}^{2}(N)$ from $\Omega_{\mathrm{DE}}(N)$ since
@@
$E_{\mathrm{DE}}^{2}$ $=\exp(-3N)\left(\frac{\Omega_{\mathrm{m}0}}{\Omega_{\mathrm{m}}}\right)=\exp(-3N)\left(\frac{1-\Omega_{\mathrm{DE}0}}{1-\Omega_{\mathrm{DE}}}\right)$ (14b)
@@
$q=-\frac{a\ddot{a}}{\dot{a}^{2}}=-1-\frac{H^{\prime}}{H},$ (15)
@@
In the following sections we will discuss quintessence for which $w_{\mathrm{DE}}=w_{\phi}$. As we will see, with the exception of ‘scaling freezing quintessence’ models, discussed in Appendix B, the evolution of the universe is described by ‘attractor’ solutions for which $\lim_{N\rightarrow-\infty}w_{\phi}=-1$ for thawing quintessence and $\lim_{N\rightarrow-\infty}w_{\phi}=w_{\infty}=\mathrm{constant}$, $-1<w_{\infty}<0$ for tracking quintessence for asymptotically inverse power law potentials. Thus, the evolution toward the past of these solutions is described by the $\Lambda$CDM and $w$CDM models, respectively. Let us therefore recall the properties of these models.
@@ 2.4 Past DE series expansions
Equation (14a) can be written as
$\Omega_{\mathrm{DE}}^{\prime}=-3w_{\mathrm{DE}}\Omega_{\mathrm{DE}}(1-\Omega_{\mathrm{DE}}).$ (20)
As we will see, thawing quintessence and tracking quintessence are associated with $\lim_{N\rightarrow-\infty}w_{\phi}=w_{\infty}=-1$ and $\lim_{N\rightarrow-\infty}w_{\phi}=w_{\infty}=\mathrm{constant}\in(-1,0)$, respectively. To derive some common features of these two types of quintessence, we therefore assume that
$w_{\infty}=\lim_{N\rightarrow-\infty}w_{\mathrm{DE}}=\mathrm{constant}\in[-1,0),$ (21)
which in combination with (20) yields
$\lim_{N\rightarrow-\infty}\Omega_{\mathrm{DE}}=0.$ (22)
Thus, $\lim_{N\rightarrow-\infty}\Omega_{\mathrm{m}}=1$, i.e. the asymptotic past is matter dominated. Equations. (20), (21) and (22) lead to the approximation $\Omega_{\mathrm{DE}}^{\prime}\approx-3w_{\infty}\Omega_{\mathrm{DE}}$ when $N\rightarrow-\infty$ and hence
$\Omega_{\mathrm{DE}}\approx Ce^{-3w_{\infty}N}.$ (23)
Motivated by this, let us now consider $w_{\mathrm{DE}}$ and $\Omega_{\mathrm{DE}}$ and perform series expansions in
@@
where $T\rightarrow 0$ when $N\rightarrow-\infty$. If one starts with a series expansion in $w_{\mathrm{DE}}$ and uses (7) and (13) one obtains an expression for $\Omega_{\mathrm{DE}}$ that subsequently can be series expanded in $T$; on the other hand, if one starts with a series expansion in $\Omega_{\mathrm{DE}}$ and inserts this into (14a) one obtains an expression for $w_{\mathrm{DE}}$ that one then can series expand in $T$. These two procedures result in the same series expansions, which can be written as [^4] The given expressions can be generalized by writing $w_{\mathrm{DE}}=w_{\infty}\left[1+\alpha T+\sigma T^{2}+\dots\right]$, which subsequently can be used to compute the series expansions for $\Omega_{\mathrm{DE}}$ and, e.g., $E_{\mathrm{DE}}^{2}$. The reason we have written $\alpha$ as $-(\gamma-1)$ and $\sigma$ as $(\gamma-1)\beta$ is that this simplifies and is compatible with our quintessence approximations.
@@ 2.5 DE Padé approximants and present initial data
For brevity we will restrict our considerations by following the most common approach in the literature, which is to use $w_{\mathrm{DE}}$ as the starting point. Next we will transform the series expansion for $w_{\mathrm{DE}}$ in $T$ to Padé approximants for two reasons:
  • (i)
    To obtain exact DE models that naturally generalize the $\Lambda$CDM and $w$CDM models and at the same time form a unifying DE context for some of the structures that thawing and tracking quintessence share. Moreover, the Padé approximants create a link between the past and the present day initial data $\rho_{\mathrm{m}0}$ and $H_{0}$ via $\Omega_{\mathrm{DE}0}=1-\Omega_{\mathrm{m}0}=1-\rho_{\mathrm{m}0}/3H_{0}^{2}$. This section thereby forms a unifying starting point for the quintessence approximations in the next section.
  • (ii)
    The main point of the paper: To obtain sufficiently good analytic approximations to the solutions of the quintessence field equations so that numerical investigations of observational issues for thawing and tracking quintessence become unnecessary! This is accomplished by transforming the series expansions that approximate solutions to the quintessence equations in the next section toward the past to Padé approximants that also are accurate at the present time. [^5] It is quite common to use Padé approximants to improve convergence in a wide range of contexts, e.g. to obtain accurate approximations for solutions to differential equations; for a brief review, references, and applications in a cosmological setting, see, e.g., [34].
Our first and second order expansions in $T$ for $w_{\mathrm{DE}}$ yield the following Padé approximants for $w_{\mathrm{DE}}$, respectively, which thereby serve as our DE models:
$w_{\mathrm{DE}}$ $=[0/1]_{w_{\mathrm{DE}}}(T)=\frac{w_{\infty}}{1+(\gamma-1)T},$ (26a)
$w_{\mathrm{DE}}$ $=[1/1]_{w_{\mathrm{DE}}}(T)=w_{\infty}\left(\frac{1+[\beta-(\gamma-1)]T}{1+\beta T}\right)=w_{\infty}\left(1-\frac{(\gamma-1)T}{1+\beta T}\right),$ (26b)
where we recall that $T=T_{0}e^{-3w_{\infty}N}$. Note that setting $\beta=\gamma-1$ in $w_{\mathrm{DE}}=[1/1]_{w_{\mathrm{DE}}}(T)$ yields $w_{\mathrm{DE}}=[0/1]_{w_{\mathrm{DE}}}(T)$ (and similarly for the expressions below for $\Omega_{\mathrm{DE}}$ and $E_{\mathrm{DE}}$ that follow from $w_{\mathrm{DE}}=[1/1]_{w_{\mathrm{DE}}}(T)$). Furthermore, setting $\gamma=1$ results in $w_{\mathrm{DE}}=w_{\infty}$ and hence $\Lambda$CDM when $w_{\infty}=-1$ and $w$CDM when $-1<w_{\infty}<0$.
To connect with initial data at the present time we can set $N=0$ and solve for $T(0)=T_{0}$ in (26) in terms of the left hand side $w_{\mathrm{DE}}(0)=w_{\mathrm{DE}0}$. Since results obtained from (26a) can be obtained from the corresponding ones obtained from (26b) by setting $\beta=\gamma-1$ we now focus on the latter. It follows from (26b) and (7) that the $[1/1]_{w_{\mathrm{DE}}}(T)$ approximant yields
$T_{0}$ $=\frac{\frac{w_{\infty}}{w_{\mathrm{DE}0}}-1}{\beta-\frac{w_{\infty}}{w_{\mathrm{DE}0}}\left[\beta-(\gamma-1)\right]},$ (27a)
$w_{\mathrm{DE}}$ $=w_{\infty}\left(1-\frac{(\gamma-1)\left(\frac{w_{\infty}}{w_{\mathrm{DE}0}}-1\right)e^{-3w_{\infty}N}}{\beta-\frac{w_{\infty}}{w_{\mathrm{DE}0}}\left[\beta-(\gamma-1)\right]+\beta\left(\frac{w_{\infty}}{w_{\mathrm{DE}0}}-1\right)e^{-3w_{\infty}N}}\right),$ (27b)
$f$ $=e^{-3w_{\infty}N}\left(\frac{\left(\gamma-1\right)\frac{w_{\infty}}{w_{\mathrm{DE}0}}}{\beta-\frac{w_{\infty}}{w_{\mathrm{DE}0}}\left[\beta-(\gamma-1)\right]+\beta\left(\frac{w_{\infty}}{w_{\mathrm{DE}0}}-1\right)e^{-3w_{\infty}N}}\right)^{\frac{\gamma-1}{\beta}},$ (27c)
where $f$ yields $\Omega_{\mathrm{DE}}$, $\Omega_{\mathrm{m}}$ and $E_{\mathrm{DE}}$ via (13) and (12). As pointed out previously, the corresponding expressions obtained from the $[0/1]_{w_{\mathrm{DE}}}(T)$ approximant are conveniently obtained from the results in (27) by setting $\beta=\gamma-1$.
@@
We note, however, that $w_{\mathrm{DE}0}$ is not an observable. It is determined by $\Omega_{\mathrm{DE}0}$ and $\left.\Omega_{\mathrm{DE}}^{\prime}\right|_{N=0}$ (or $\left.\frac{\Omega_{\mathrm{DE}}}{dz}\right|_{z=0})$) via eq. (14a), which is not an easy observational task. For this reason we will instead relate $T_{0}$ to $\Omega_{\mathrm{DE}0}$ and thereby obtain a direct connection to the observables $H_{0}$ and $\rho_{\mathrm{m}0}$ since $1-\Omega_{\mathrm{DE}0}=\Omega_{\mathrm{m}0}=\rho_{\mathrm{m}0}/3H_{0}^{2}$. We therefore consider the $[1/1]_{\Omega_{\mathrm{DE}}}(T)$ Padé approximant for $\Omega_{\mathrm{DE}}$ (use the expression (25b) up to second order in $T$):
$\Omega_{\mathrm{DE}}=[1/1]_{\Omega_{\mathrm{DE}}}(T)=\frac{T}{1+\gamma T},$ (28)
@@
from which it follows that $\Omega_{\mathrm{DE}0}=T_{0}/[1+\gamma T_{0}]$, which results in